# appAdditional References Appendix Ensuring Medical AI Safety: Explainable AI-Driven Detection and Mitigation of Spurious Model Behavior and Associated Data

### v1

@@ Abstract
Deep neural networks are increasingly employed in high-stakes medical applications, despite their tendency for shortcut learning in the presence of spurious correlations, which can have potentially fatal consequences in practice. Detecting and mitigating shortcut behavior is a challenging task that often requires significant labeling efforts from domain experts. To alleviate this problem, we introduce a semi-automated framework for the identification of spurious behavior from both data and model perspective by leveraging insights from eXplainable Artificial Intelligence (XAI). This allows the retrieval of spurious data points and the detection of model circuits that encode the associated prediction rules. Moreover, we demonstrate how these shortcut encodings can be used for XAI-based sample- and pixel-level data annotation, providing valuable information for bias mitigation methods to unlearn the undesired shortcut behavior. We show the applicability of our framework using four medical datasets across two modalities, featuring controlled and real-world spurious correlations caused by data artifacts. We successfully identify and mitigate these biases in VGG16, ResNet50, and contemporary Vision Transformer models, ultimately increasing their robustness and applicability for real-world medical tasks.
@@ 1 Introduction
In the past decade, Artificial Intelligence (AI) models have become ubiquitous in medical applications, often outperforming human experts in tasks like melanoma detection [14] and the prediction of cardiovascular diseases from electro cardiogram (ECG) data [68]. However, the non-transparent nature of Deep Neural Network (DNN) predictions poses challenges in safety-critical contexts, as their reasoning remains obscure to both clinicians and model developers. This opacity is concerning, particularly since DNNs are prone to exploit spurious correlations in the training data. This can lead to shortcut learning [39], where models rely on (medically) irrelevant features, yet correlating with the target label. Such shortcuts are not limited to protected attributes like gender or ethnicity, but include various confounders in the training data, such as objects (e.g., rulers or hair), color shifts, or watermarks added by scanning devices. A well-known example are band-aids in dermoscopic images for melanoma detection dominantly occurring beside benign lesions, causing AI models to associate the presence of band-aids with benign lesions with potentially severe consequences in practice. Similarly, DNNs trained to detect pneumonia from radiographs have predicted the hospital system used for the scan, as prevalence varied across hospitals in the study [78]. Moreover, confounding shortcuts were learned over intended signals from computed tomography scans for COVID-19 detection [23].
@@
The field of XAI sheds light onto the previously non-transparent prediction behavior of DNNs, providing insights into their internal reasoning. While traditional local XAI methods focus on feature importance for individual predictions, global XAI approaches and mechanistic interpretability aim to understand overall model behavior by explaining the roles of internal representations and encoded features [1, 34, 79]. Recently, these insights have been utilized to systematically uncover model weaknesses like shortcut behavior. Current methods either detect outlier explanations for biased data samples [48, 3, 27] or outlier model concept representations [16, 53, 59]. Beyond revealing model weaknesses, XAI is also capable of identifying data-related issues causing the undesired behavior. Most related to our work, the iterative Reveal2Revise framework [59] consists of the four steps (a) bias revealing, (b) bias modeling, i.e., learning accurate bias representations, (c) model revision to unlearn shortcuts, and (d) re-evaluation. However, while it primarily focuses on model revision, it provides limited insights into bias revelation and modeling steps. This can lead to imperfect bias representations, limiting the effectiveness of the mitigation step. To fill this gap, we enhance the Reveal2Revise framework by providing a robust methodological foundation for the bias revelation and modeling steps to (semi-)automatically annotate and enrich datasets, leveraging insights from XAI, as illustrated in Fig. 1. Specifically, we discuss bias identification from both data and model perspectives, ( ) enabling expert-guided validation of model behavior. Moreover, ( ) we learn a model-internal bias representation, referred to as bias model, from an initial set of biased samples using Concept Activation Vectors (CAVs). We then ( ) introduce the retrieval of biased data samples using the learned bias model, ( ) enabling its iterative improvement. Furthermore, ( ) using the refined bias model for spatial bias localization, we enrich the dataset with extracted sample- and pixel-level annotations, which are used in the bias mitigation and evaluation steps of the Reveal2Revise framework to improve the generalization capabilities of AI models.
@@
We demonstrate the applicability of our framework on four medical tasks across two modalities and provide annotations for the detected spurious concepts and data points111Both code and annotations are provided here: https://github.com/frederikpahde/medical-ai-safety. This includes image-based melanoma detection, the identification of gastrointestinal abnormalities, cardiomegaly prediction from chest radiographs, and cardiovascular disease prediction from ECG data using VGG, ResNet, and Vision Transformer (ViT) models. Utilizing our framework, we identify and mitigate both intrinsic and artificially introduced biases across all datasets, and demonstrate the data annotation capabilities using concept-based bias representations, minimizing the need for human supervision to improve the validity and robustness of AI models.
@@
The paper is structured as follows: After summarizing related work in Sec. 2, we discuss how biases and concepts in general are modeled in DNNs (Sec. 3.1), and leverage these representations for the (iterative) retrieval of biased samples (Sec. 3.2) and spatial bias localization (Sec. 3.3). We further discuss how (outlier) concepts and data points can be detected in the first place (Sec. 4), briefly describe bias mitigation methods (Sec. 5), and demonstrate the applicability of our framework in experiments (Sec. 6). Lastly, we discuss limitations and conclusions in Secs. 7 and 8, respectively.
2 Related Work
@@
Existing XAI-based bias identification approaches primarily focus on shortcut detection in pre-trained models on benchmark datasets like ImageNet [24]. One line of research aims to identify samples with outlier model behavior, measured via local attribution scores in input [48] or latent [3, 27] space. Other methods seek to pinpoint spurious model representations like neurons or circuits directly. For instance, Singla and Feizi [66] employ human annotators to label neurons as valid or spurious using feature visualization techniques, leveraging this information to detect spurious samples. To minimize manual labeling, they consider only a subset of neurons based on their mutual information with model failures. In contrast to individual neurons, Neuhaus et al. [53] define class-wise neural Principal Component Analysis (PCA) to identify meaningful latent directions for each class, and leverage human annotators to label a pre-selected subset as spurious. With this approach they identify spurious features, such as the presence of bird feeders impacting the prediction of hummingbirds. Another recent bias identification method is Data-agnOstic Representation Analysis (DORA) [16], which generates samples maximizing the neurons activations to represent concepts and defines a distance measure based on how neurons respond to representations of other neurons to identify outlier concepts. This has resulted in the identification of neurons extracting spurious concepts, such as watermarks or background features. In a medical context, existing works study shortcut behavior related to sensitive attributes like gender or age by comparing the model performance in different sub-popluations [15], manually annotate the dataset for spurious features [9], or define heuristics to automate the detection of specific artifacts [62]. In contrast, our work presents a generic framework for the XAI-based identification of spurious correlations without prior knowledge and the (semi-)automatic generation of sample- and pixel-level annotations for spurious features. Closest to our work is Reveal2Revise [59], an iterative model correction framework with the steps (1) reveal for bias identification, (2) bias modeling, (3) revise for bias mitigation, and (4) (re-)evaluation. While the Reveal2Revise framework emphasizes bias mitigation (step 3), our work extends and examines the XAI-based bias identification and modeling (steps 1 and 2) in greater detail, as well as its applicability for data annotation to reduce manual labeling efforts.
@@ 3 From Bias Modeling to (Semi-)Automated Data Annotation
In recent years, the XAI community has shifted its focus from local to global explanations to better understand overall model behavior. This line of research, known as mechanistic interpretability, aims to interpret internal representations in terms of human-understandable concepts, encoded as individual neurons, model circuits, or directions in latent space. Understanding model internals enables the identification of model substructures that encode biases stemming from data artifacts (step ( ) in Fig. 1). In this section, we discuss how biases are encoded within DNNs, and how this understanding can be utilized for data annotation tasks, such as the detection of biased samples and (spatial) bias localization. Note that while we assume knowledge on the existence of biases in this section, the identification thereof (step ( ) in Fig. 1) is addressed in Sec. 4.
@@
3.1 Bias Modeling from an Explainable AI Perspective
@@
Traditionally, it has been assumed that neurons in robust models encode human-aligned concepts, particularly at layers close to the model head [55, 61, 32, 6]. Hence, there might exist a neuron acting as feature extractor for a biased concept. Various feature visualization approaches aim to globally explain the concept represented by a neuron by identifying inputs that maximally trigger the neuron. Whereas one line of work generates inputs that maximally activate the selected neuron [33, 55, 35], other approaches select natural images from a reference dataset, e.g., the training set. Specifically, while Activation Maximization (ActMax) [69, 11] selects samples that maximally activate a given neuron, Relevance Maximization (RelMax) [1] selects samples for which the neuron is maximally relevant for the classification task, as computed by a local explainability methods. In contrast to activations, the relevance scores are directly linked to the model prediction, indicating the neuron’s impact on a specified target label. However, limitations of the mapping of concepts to individual neurons are redundancy, i.e., multiple neurons representing the same concept [25], and polysemanticity [38, 56, 31, 29], i.e., neurons reacting to multiple, seemingly unrelated concepts. Recent works aim to overcome these challenges by disentangling learned concepts via Sparse Autoencoders (SAEs) [45, 13]. Assuming there are more concepts than neurons, SAEs leverage sparse dictionary learning to find an overcomplete feature basis, allowing the usage of encoder neurons as monosemantic concept representation.
@@
Supervised Concept Estimation: When concept labels indicating the presence of a bias are known, concept directions can be estimated via Linear Probing [2, 7]. As such, CAVs, introduced for concept sensitivity testing [47], were originally computed as the weight vector from a linear classifier like a linear SVM, distinguishing latent activations on layer of samples with the concept from activations of samples without it. However, recent work has shown that CAVs derived from linear classifiers can be influenced by distractors in the data, leading to inaccurate estimates of the concept signal direction [58]. The authors argue that maximizing class separability is the wrong optimization target and propose Pattern-CAVs, as detailed in Appendix A.1. CAVs can be computed on latent activations of arbitrary input shape, such as 3D representations of shape , representing channel, height, and width dimensions, as well as aggregated representations. This includes translation-invariant 1D or channel-invariant 2D representations by using max- or average pooling over spatial or channel dimensions [50]. In this work, we use 1D-CAVs with max-pooling over spatial dimensions. Beyond linear concept representations, Concept Activation Regions have been proposed [21], allowing for non-linear concept representations through a radial kernel function.
@@
Unsupervised Concept Discovery: When concept labels are not available, concepts can be discovered in unsupervised fashion. An early approach, Automatic Concept Extraction (ACE) [40], extracts concepts by segmenting images into regions and clustering similar regions to identify potential visual concepts. In contrast, recent approaches leverage matrix decomposition methods, such as PCA, Singular Value Decomposition (SVD), or Non-negative Matrix Factorization (NMF), on latent activations to identify meaningful directions in the model’s latent space [79, 34, 41, 53]. This leads to two matrices, with one matrix reinterpreted as the concept basis, i.e., each row can be considered as CAV, and the other matrix as the activations within that new basis [36]. Another work proposes the discovery of linear subspaces as concept representations [72]. However, unsupervised concept discovery for representing data artifacts has two drawbacks: First, while no labeling efforts are needed to find concept labels, manual inspection is required to determine which direction(s) accurately represent the targeted concept, i.e., the artifact to be modeled. Second, matrix factorization approaches yield statistical groupings without guidance without a guarantee for the existence of a direction representing the targeted artifact concept.
3.2 Biased Sample Retrieval
@@
With a precise bias representation via CAV or neuron , the detection of artifactual samples can be further automated (step ( ) in Fig. 1). Specifically, all samples from a dataset can be ranked by their similarity to the artifact representation, computed for example via cosine similarity, and presented to human annotators in that order, significantly supporting them in detecting artifact samples. Therefore, the data annotation process for concept representations based on a single neuron , as for example suggested by Singla and Feizi [66], is similar to global XAI methods explaining individual neurons via ActMax or RelMax, which retrieve reference samples with maximal activation or relevance for the given neuron. However, the limitations discussed above, namely redundancy and polysemanticity, affect the concept detection capabilities of individual neurons. To address this, extending the artifact representation to linear directions in latent space via CAVs is a viable solution. Specifically, given CAV and sample , we can compute a bias score by projecting latent activations for layer onto the CAV:
@@
Alternatively, inspired by RelMax, we can compute bias scores using relevance scores instead of activations, as outlined in Appendix A.2. Since relevance scores are computed class-specifically, this approach allows distinguishing concepts that are artifactual for certain classes but valid for others. An illustration is provided in Fig. 2 (left), where a CAV has been trained to distinguish between known artifact () and non-artifact samples (). This concept representation can subsequently be used to compute concept scores for unknown samples (), supporting annotators in the detection of further artifact samples.
@@
Concept representations can be refined iteratively by correcting labeling errors in the data (step ( ) in Fig. 1). Specifically, non-artifactual and unknown samples with high bias scores are subject to manual inspection to improve the label quality and concept representation, as shown in Fig. 3. Starting with ( ) a small set of bias samples obtained from bias identification methods, ( ) an initial CAV is fitted. Next, ( ) manual inspection of samples with high bias scores improves the label quality. The updated labels are used to ( ) iteratively refine the CAV, ( ) resulting in a set of annotated bias samples.
3.3 Spatial Bias Localization
Beyond detecting artifact samples, XAI insights can further reduce human labeling efforts by automating the spatial localization of biased (and localizable) concepts within these samples (step ( ) in Fig. 1). We assume the existence of a bias representation via CAV or neuron . The latter can be represented as CAV by setting all values to zero, except for the one for neuron , which is set to one. The targeted concept can then be localized in input space using local attribution methods, such as Layer-wise Relevance Propagation (LRP) [4]. Using singular neurons as concept representation, Singla and Feizi [66] leverage Class Activation Maps [80] to visualize the feature map for the given neuron in input space and Achtibat et al. [1] mask latent relevances for local attribution methods to compute neuron-specific input heatmaps. This approach can be extended to biases represented as direction in latent space, e.g., using PCA [53] or CAVs [3, 59, 22]. Specifically, given the concept direction and latent activations , we can utilize local explanation approaches like LRP [4] to explain the prediction of the function producing the bias score as defined in Eq. 1. This is equivalent to a local attribution method applied to sample , with latent relevance scores initialized as

### v2

# appAdditional References Appendix Ensuring Medical AI Safety: Interpretability-Driven Detection and Mitigation of Spurious Model Behavior and Associated Data

@@ Abstract
Deep neural networks are increasingly employed in high-stakes medical applications, despite their tendency for shortcut learning in the presence of spurious correlations, which can have potentially fatal consequences in practice. Whereas a multitude of works address either the detection or mitigation of such shortcut behavior in isolation, the Reveal2Revise approach provides a comprehensive bias mitigation framework combining these steps. However, effectively addressing these biases often requires substantial labeling efforts from domain experts. In this work, we review the steps of the Reveal2Revise framework and enhance it with semi-automated interpretability-based bias annotation capabilities. This includes methods for the sample- and feature-level bias annotation, providing valuable information for bias mitigation methods to unlearn the undesired shortcut behavior. We show the applicability of the framework using four medical datasets across two modalities, featuring controlled and real-world spurious correlations caused by data artifacts. We successfully identify and mitigate these biases in VGG16, ResNet50, and contemporary Vision Transformer models, ultimately increasing their robustness and applicability for real-world medical tasks. Our code is available at https://github.com/frederikpahde/medical-ai-safety.
@@ 1 Introduction
In the past decade, Machine Learning (ML) models have become ubiquitous in medical applications, often outperforming human experts in tasks like melanoma detection [15] and the prediction of cardiovascular diseases from electro cardiogram (ECG) data [69]. However, the non-transparent nature of Deep Neural Network (DNN) predictions poses challenges in safety-critical contexts, as their reasoning remains obscure to both clinicians and model developers. This opacity is concerning, particularly since DNNs are prone to exploit spurious correlations in the training data. This can lead to shortcut learning [39], where models rely on (medically) irrelevant features, yet correlating with the target label. Such shortcuts are not limited to protected attributes like gender or ethnicity, but include various confounders in the training data, such as objects (e.g., rulers or hair), color shifts, or watermarks added by scanning devices. A well-known example are band-aids in dermoscopic images for melanoma detection dominantly occurring alongside benign lesions, causing ML models to associate the presence of band-aids with benign lesions with potentially severe consequences in practice. Similarly, DNNs trained to detect pneumonia from radiographs have been known to predict the hospital system used for the scan, as the prevalence varied across hospitals in the study [79]. Moreover, confounding shortcuts were learned over intended signals from computed tomography scans for COVID-19 detection [24].
@@
The field of XAI sheds light onto the previously non-transparent prediction behavior of DNNs, providing insights into their internal reasoning. While traditional local XAI methods focus on feature importance for individual predictions, global XAI approaches aim to understand overall model behavior by explaining the roles of internal representations and encoded features [1, 34, 80]. Recently, these insights have been utilized to uncover model weaknesses like shortcut behavior. Current methods either detect outlier explanations for biased data samples [48, 3, 28] or outlier model concept representations [17, 52, 57]. The purpose of this work is to review interpretability-driven shortcut detection and bias mitigation methods. Specifically, we focus on comprehensive frameworks such as Reveal2Revise [57], which consists of the four steps (a) bias revealing, (b) bias modeling, i.e., learning accurate bias representations, (c) model revision to unlearn shortcuts, and (d) re-evaluation. While we review related work in the first three steps, we put special attention on the bias detection and modeling steps. We further extend the framework with bias annotation capabilities to (semi-)automatically annotate and enrich datasets, leveraging insights from XAI, as shown in Fig. 1 (see Appendix A.1 for an algorithmic overview). Specifically, we discuss bias identification from both data and model perspectives, enabling expert-guided validation of model behavior111Ideally, this is done by an interdisciplinary group consisting of model developers and domain experts. ( ). Moreover, we learn a model-internal bias representation, referred to as bias model, from an initial set of biased samples using Concept Activation Vectors (CAVs) ( ). We then introduce the retrieval of biased data samples using the learned bias model ( ), enabling its iterative improvement ( ). Further using the refined bias model for spatial bias localization ( ), we enrich the dataset with extracted sample- and feature-level bias annotations, which are used in the bias mitigation and evaluation steps of the Reveal2Revise framework to improve the generalization capabilities of ML models.
@@
We demonstrate the applicability of the extended framework on four medical tasks across two modalities and provide annotations for the detected spurious concepts and data points222Both code and annotations are provided here: https://github.com/frederikpahde/medical-ai-safety. This includes image-based melanoma detection, the identification of gastrointestinal abnormalities, cardiomegaly prediction from chest radiographs, and cardiovascular disease prediction from ECG data using VGG, ResNet, and Vision Transformer (ViT) models. Utilizing Reveal2Revise, we identify and mitigate both intrinsic and artificially introduced biases across all datasets, and demonstrate the data annotation capabilities using concept-based bias representations, minimizing the need for human supervision to improve the validity and robustness of ML models.
@@
The paper is structured as follows: First, we summarize related work on comprehensive bias detection and mitigation frameworks in Sec. 2. We then provide background on how (biased) concepts are encoded within DNNs (Sec. 3.1). Our main contribution lies in utilizing these representations for sample- and feature-level bias annotation, specifically the (iterative) retrieval of biased samples (Sec. 3.2) and spatial bias localization (Sec. 3.3).We further discuss and categorize existing methods to detect (outlier) concepts and data points (Sec. 4), briefly describe existing bias mitigation methods (Sec. 5), and demonstrate the applicability of the extended Reveal2Revise framework in experiments (Sec. 6). Lastly, we discuss limitations and conclusions in Secs. 7 and 8, respectively.
2 Related Bias Detection and Mitigation Frameworks
@@
Many existing interpretability-based bias identification methods focus on structured data, which allows for clear concept specifications, particularly in tabular datasets [7, 60]. In contrast, works involving unstructured data such as images primarily concentrate on shortcut detection in pre-trained models using generic benchmark datasets like ImageNet [25]. One research directions aims to identify samples with outlier model behavior, measured via local attribution scores in input [48] or latent [3, 28] space. Other methods seek to pinpoint spurious model representations like neurons [67] or latent directions [52], and modify them for bias mitigation. n a medical context, existing works study shortcut behavior related to sensitive attributes like gender or age by comparing the model performance in different sub-popluations [16], manually annotate the dataset for spurious features [10], or define heuristics to automate the detection of specific artifacts [62]. In contrast, our work emphasizes generic and comprehensive methods for bias detection and mitigation, such as Reveal2Revise. This iterative model correction framework consists of four steps: (1) Reveal for bias identification, (2) bias modeling, (3) Revise for bias mitigation, and (4) (re-)evaluation (see Appendix A.1). Reveal2Revise enables the interpretability-based identification and mitigation of spurious correlations without requiring prior knowledge. While the original Reveal2Revise work [57] emphasizes bias mitigation (step 3), our paper reviews the interpretability-based bias identification and modeling (steps 1 and 2) in greater detail. We further extend the framework with bias annotation techniques, enabling the (semi-)automated generation of sample- and feature-level bias annotations. This significantly reduces the manual labeling efforts necessary for obtaining annotations required for bias mitigation methods. Similar to Reveal2Revise, Schramowski et al. [64] provide a comprehensive framework that allows domain experts to interact with local explanations to align ML models with scientifically expected behavior.
@@ 3 From Bias Modeling to (Semi-)Automated Data Annotation
Both medically valid and biased concepts manifest in specific model components, such as individual neurons, model circuits, or directions in latent space. In this section, we review methods from global XAI and mechanistic interpretability that aim to interpret model substructures, specifically those encoding biases stemming from data artifacts (step ( ) in Fig. 1). We further utilize these model insights to extend the Reveal2Revise frameworks with bias annotation capabilities, including the detection of biased samples and (spatial) bias localization. Note that while we assume knowledge on the existence of biases in this section, the identification thereof (step ( ) in Fig. 1) is addressed in Sec. 4.
@@
3.1 Background: Concept Representations in Neural Networks
@@
A common assumption is that neurons in robust models encode human-aligned concepts, particularly at layers close to the model head [54, 61, 6]. Hence, there might exist a neuron acting as feature extractor for a biased concept. Various feature visualization approaches aim to globally explain the concept represented by a neuron by identifying inputs that maximally trigger the neuron. Whereas one line of work generates inputs that maximally activate the selected neuron [33, 54, 35], other approaches select natural images from a reference dataset, e.g., the training set. Specifically, while Activation Maximization (ActMax) [70, 12] selects samples that maximally activate a given neuron, Relevance Maximization (RelMax) [1] selects samples for which the neuron is maximally relevant for the classification task, as computed by a local explainability methods. In contrast to activations, the relevance scores are directly linked to the model prediction, indicating the neuron’s impact on a specified target label. However, limitations of the mapping of concepts to individual neurons are redundancy, i.e., multiple neurons representing the same concept [26], and polysemanticity [38, 55, 32, 30], i.e., neurons reacting to multiple, seemingly unrelated concepts. Recent works aim to overcome these challenges by disentangling learned concepts via Sparse Autoencoders (SAEs) [45, 14]. Assuming there are more concepts than neurons, SAEs leverage sparse dictionary learning to find an overcomplete feature basis, allowing the usage of encoder neurons as monosemantic concept representation.
@@
Supervised Concept Estimation: When concept labels indicating the presence of a bias are known, concept directions can be estimated via Linear Probing [2, 8]. As such, CAVs, introduced for concept sensitivity testing [47], were originally computed as the weight vector from a linear classifier like a linear SVM, distinguishing latent activations on layer of samples with the concept from activations of samples without it. However, recent work has shown that CAVs derived from linear classifiers can be influenced by distractors in the data, leading to inaccurate estimates of the concept signal direction [58]. Instead of maximizing class separability, which is considered as the wrong optimization target, the work proposes Pattern-CAVs as robust alternative, as detailed in Appendix A.2. CAVs can be computed on latent activations of arbitrary input shape, such as 3D representations in , with and representing the height and width dimensions, respectively, as well as aggregated representations. This includes translation-invariant 1D or channel-invariant 2D representations by using max- or average pooling over spatial or channel dimensions [50]. In this work, we use 1D-CAVs with max-pooling over spatial dimensions. Beyond linear concept representations, Concept Activation Regions have been proposed [22], allowing for non-linear concept representations through a radial kernel function.
@@
Unsupervised Concept Discovery: When concept labels are not available, concepts can be discovered in unsupervised fashion. An early approach, Automatic Concept Extraction (ACE) [40], extracts concepts by segmenting images into regions and clustering similar regions to identify potential visual concepts. In contrast, more recent approaches leverage matrix decomposition methods, such as Principal Component Analysis (PCA), Singular Value Decomposition (SVD), or Non-negative Matrix Factorization (NMF), on latent activations to identify meaningful directions in the model’s latent space [80, 34, 41, 52]. This leads to two matrices, with one matrix reinterpreted as the concept basis, i.e., each row can be considered as CAV, and the other matrix as the activations within that new basis [36]. Another work proposes the discovery of linear subspaces as concept representations [73]. However, unsupervised concept discovery for representing data artifacts has two drawbacks: First, while no labeling efforts are needed to find concept labels, manual inspection is required to determine which direction(s) accurately represent the desired concept, i.e., the artifact to be modeled. Second, matrix factorization approaches yield statistical groupings without guidance, such that there is no guarantee for the existence of a direction representing the targeted artifact concept.
3.2 Sample-Level Bias Annotation via Data Retrieval
@@
With a precise bias representation via CAV or neuron , the detection of artifactual samples can be further automated (step ( ) in Fig. 1). Specifically, all samples from a dataset can be ranked by their similarity to the artifact representation, computed for example via cosine similarity, and presented to human annotators in that order, significantly supporting them in detecting artifact samples. Therefore, the data annotation process for concept representations based on a single neuron , as for example suggested by Singla and Feizi [67], is similar to global XAI methods explaining individual neurons via ActMax or RelMax, which retrieve reference samples with maximal activation or relevance for the given neuron. However, the limitations discussed above, namely redundancy and polysemanticity, affect the concept detection capabilities of individual neurons. To address this, we propose the extension of the artifact representation to linear directions in latent space via CAVs. Specifically, given CAV and sample , we can compute a bias score by projecting latent activations for layer onto the CAV:
@@
High bias scores indicate a higher likelihood for the presence of the modeled bias. Alternatively, inspired by RelMax, we can compute bias scores using relevance scores instead of activations, as outlined in Appendix A.3. Since relevance scores are computed class-specifically, this approach allows distinguishing concepts that are artifactual for certain classes but valid for others. An illustration is provided in Fig. 2 (left), where a CAV has been trained to distinguish between known artifact () and non-artifact samples (). This concept representation can subsequently be used to compute concept scores for unknown samples (), supporting annotators in the detection of further artifact samples.
@@
Concept representations can be refined iteratively by correcting labeling errors in the data (step ( ) in Fig. 1). Specifically, non-artifactual and unknown samples with high bias scores are subject to manual inspection to improve the label quality and concept representation, as shown in Fig. 3. Starting with a small set of bias samples obtained from bias identification methods ( ), an initial CAV is fitted ( ). Next, manual inspection of samples with high bias scores improves the label quality ( ). The updated labels are used to iteratively refine the CAV ( ), resulting in a set of annotated bias samples ( ).
3.3 Feature-Level Bias Annotation via Spatial Localization
Beyond detecting artifact samples, XAI insights can further reduce human labeling efforts by automating the spatial localization of biased (and localizable) concepts within these samples (step ( ) in Fig. 1). We assume the existence of a bias representation via CAV or neuron . The latter can be represented as CAV by setting all values to zero, except for the one for neuron , which is set to one. The targeted concept can then be localized in input space using local attribution methods, such as Layer-wise Relevance Propagation (LRP) [4]. Assuming singular neurons as concept representation, Singla and Feizi [67] leverage Class Activation Maps [81] to visualize the feature map for the given channel in input space and Achtibat et al. [1] mask latent relevances for local attribution methods to compute channel-specific input heatmaps. This approach can be extended to biases represented as direction in latent space, e.g., using PCA [52] or CAVs [3, 57, 23]. Specifically, given the concept direction and latent activations , we can utilize local explanation approaches like LRP [4] to explain the prediction of the function producing the bias score as defined in Eq. 1. This is equivalent to a local attribution method applied to sample , with latent relevance scores initialized as